From GlueXWiki
Jump to: navigation, search

February 28, 2013 FDC meeting


  1. Production Construction Tracking (Dave)
    • Spare parts
    • Spare wire plane testing
    • Package #2 status
  2. Rebaselining update [1](Lubomir)
  3. Engineering (Bill)
    • Cooling system test
  4. Electronics (Chris, Nick)
  5. Tests at EEL126 FDC E-log (Beni)
  6. Other


Participants: Eugene, Dave, Chris, Nick, Simon, Beni, and Lubomir.


- Dave for the spare parts: working on the last cathodes #87, 88, 89. These have been tensioned and we are working now on the daughter cards. We will build also an extra-spare cathode #90; the foils have been glued for it.

- As Casey proposed after the collaboration meeting, he replaced the six deadened wires (deadened area of 6cm diameter) of wire plane #2 with new wires. This will allow us to make 7.8cm deadened area, that is needed for the spare package. The wires are glued, will be soldered tomorrow and the tension tested.

- Spare wire plane #6 with wires strung but without HV caps, has been tested: 600/-60nA at 2100/-500V. Now Anatoly is installing the caps.

- Package #2: the grounding and the cooling tubes have been installed. Chris and Nick will come on Monday to do the HV connections. So far the package is still sitting on the mounting table.

- Fixing the field HV sector in package #3: we started discussion about possible solutions. The problem is in cell #5 and one option would be to take the top 4 cells together. We have done this before. Still, more people (Casey, Dave) are inclined to open it plane by plane (as usual) but from the bottom side. The problem is that the spacer ring is screwed to the wire frame and we don't have access to the screws from the bottom. However it should be OK with more people to take them together. Before attempting this, Lubomir will test once more this HV sector; already stared flushing package #3 with regular gas and will bring the power supply to do the test.

Rebaselining update

- We are filling the html templates that Eugene created for each activity. Where available we link documents with more details (see the file linked above and click on the 25211191 activity for ex.). Eugene wanted to have all the activities finished in this manner by the middle of the next week.


Bill was not at the meeting but some of the questions were clarified after the meeting by him.

- Due to the problems with the power in 126 (see below), Bill proposed to use a water chiller borrowed from the FEL (120VAC) both for the cooling system test and for the test of one package with full electronics in 126. The chiller requires some fixing that will be done in 5-10 days.

- Cooling system tests: Bill wants to test his ANSI model using one loop with 22 pre-amps. So far we can't find spare pre-amp back-brackets needed for the tests, but will find a solution to that. Eugene was concerned the viscosity of the water and Fluorinert are different, but Bill explained actually they are very similar. Beni: more important are the impedances and this is what this test is going to check. Bill has results from ANSI for both, water and Fluorinert.

- Cooling of one full package in 126: The water chiller is 1kW which is more than enough for one package that needs ~0.25kW. Eugene was concerned that water may leak and cause damages, and that's why we decided to use Fluorinert. Unlike inside the magnet, we will have access to all the tube connections. Also, Bill wants to test the tubing for leakage with gas and then to put water there.

- 208VAC power in 126: according to Jason Willoughby there are only 3 spaces left in the panel board which is enough for either (1) one real-3-phase 208VAC receptacle, or (2) three "1-phase" 208VAC receptacles. (1) is needed for the SMC chiller (with Fluorinert) and (2) is needed for the three VXS crates. That's why the only way to do the test with full electronics in 126 is to have three "1-phase" 208VAC receptacles and use the FEL water chiller. Bill instructed Jason to proceed with the second option.


- Decided to use "red" color to mark all the FDC cables. "Black" was taken by BCAL.

- Chris and Nick will come next week to Blue Crab to work on the HV cabling for package #2 and to populate the last two wire planes.

Tests at EEL126

- Beni continues testing the first package: taking plateau data at different thresholds in the range of 5-11Volts (on the threshold supply) and 2050-2200V on the sense wires. Will be ready with the tests by Wednesday next week. The plan is to bring the package back to Blue Crab Wednesday afternoon and eventually to move the second package (the last to be tested) to EEL126.